OpenStudy (anonymous):

Exact value of cot (315) degree

OpenStudy (mathstudent55):

What quadrant is it in?

OpenStudy (anonymous):

Quadrant IV

OpenStudy (mathstudent55):

good

OpenStudy (anonymous):

So? What do I do

OpenStudy (anonymous):

tan=perpendicular/base cot=1/tan=1/p/b=b/p 315degrees= 7pi/4 1/tan(7pi/4)

OpenStudy (mathstudent55):

What is the cotangent equal to?

OpenStudy (mathstudent55):

in terms of sin and cos, or in terms of the tan

OpenStudy (anonymous):

7pi / 4

OpenStudy (mathstudent55):